# High-cadence observations of galactic nuclei by the future two-band UV-photometry mission QUVIK

### v1

@@ 3.1 Reverberation mapping of accretion disks
Figure 5: Top panel: Simulated light curves of a reverberating accretion disk for $M_{\bullet}=10^{9}M_{\odot}$ as modeled in the FUV (175 nm), NUV1 (260 nm), and NUV (325 nm) bands. Dense light curves are plotted in faded colors, while dark points represent mock observations with cadence of 0.5 days. Lower-left panel: Discrete correlation coefficient as a function of the time lag, $\tau$, prior to bootstrapping between the FUV and NUV2 bands. Lower-right panel: Interpolated cross-correlation coefficient as a function of $\tau$ between the FUV and NUV2 bands. Within the lower panels, the blue vertical dashed line represents the peak correlation coefficient, while the pink vertical dotted line represents the centroid measured at 80 % of the peak. The black vertical dash-doted line represents the mean time delay as calculated directly from the means of the transfer functions.
@@
With the two UV bands, QUVIK has the ability to perform reverberation mapping in UV wavelengths. The AGN’s UV emission is likely due to a relatively compact region of the accretion disk near the supermassive black hole (SMBH) [68, 69, 70]. Spectra of accretion disks are more complex than a simple black body emission of typical accretion disk models (e.g. thin disk [71, 72], slim disk [73], and advection dominated accretion flows or ADAFs [74]). Magneto-hydrodynamic (MHD) simulations tend to describe the accretion flow to be composed from several gaseous components with different temperature, density, and geometric properties  [75, 76, 77, 78, 79]. The real accretion disk is likely a combination of many of these accretion flows at different size scales. While optical continuum reverberation mapping seems to support some of these accretion flows (albeit with $\sim$2 to 3 times larger accretion disk sizes [16, 80]), there are fewer studies of UV reverberation [68, 81, 82, 83]. QUVIK offers the ability to fill in this gap, with the higher cadence required to probe the smaller size scales of the inner accretion disk. In Section 4, we describe how this can be joined with other surveys at different wavelengths (optical and X-ray) to provide reverberation mapping across a wider spectrum.
@@
Here we perform simulations of AGN light curves in order to test the efficacy of QUVIK on reverberation mapping utilizing Amoeba [84]. The UV emission of the AGN is simulated as the reprocessing of a stochastic driving signal within the lamppost model, where a point-like corona irradiates the accretion disk from a position directly above the black hole [85, 86]; see also Fig. 4 for the illustration of the AGN setup that is inspired by the unified AGN model[87]. We generate light curves assuming Schwarzschild SMBH with masses of $M_{\bullet}=10^{8},10^{9}M_{\odot}$, accreting with an Eddington ratio of 10 % and placed at a redshift of 0.03. The corona height is taken to be 6 $r_{\rm{g}}$, which is equal to the inner radius of the accretion disk, and it is assumed that it radiates with an average power of 10 % of the bolometric luminosity. We then model the driving signal as a damped random walk (DRW), which has a characteristic time scale of 20 days and an asymptotic structure function value of 300 [88], emphasizing that this driving signal is not an observable quantity.
@@
From this simulation, the UV signals are computed assuming that the accretion disk thermally responds when the driving signal is absorbed. We computed this response by calculating the transfer functions of the accretion disk when viewed at 20[86]. These transfer functions encode the geometry and physics of the models used. It is known that inclined accretion disks do not have a significant change in their mean response function, but can change the skew of the transfer function. The UV signals are then the convolution of these transfer functions and the driving signal, from which we add 2 % Poisson noise and sample the dense light curves based on various observation strategies. We perform simulations for 3 UV bands: FUV, NUV1, and NUV2 using effective wavelengths of 175, 260, and 325 nm, respectively, where FUV and NUV2 correspond to the currently preferred QUVIK bands and NUV1 corresponds to the ULTRASAT band.
@@
To compute the time lag between the UV light curves, we then use two methods known as the Discrete Correlation Function (DCF) [89] and the Interpolated Cross-Correlation Function (ICCF) [90, 91]. These methods compute the correlation coefficient based on the time delay between two signals, where 1.0 represents perfect correlation. The time lag between two signals may be taken to be the time lag associated with the peak correlation coefficient, or alternatively, the centroid value at a specified level [85]. Common thresholds for the centroid value are 50 or 80 % of the peak correlation coefficient, which is found to give more reliable time lag measurements with respect to the peak [92, 93]. In doing so, we are less biased towards how sharp or skewed the transfer functions are. The uncertainties in the time lag are then computed using a bootstrapping method for the DCF and a Markov Chain Monte Carlo method for the ICCF, which improves our understanding of the time lag between the UV bands.
@@ 3.2 Stochastic nuclear transients
where $\kappa$ is of the order of unity and depends on the radial density profile of the star. Eq. (3) expresses the tidal radius conveniently in gravitational radii, $r_{\rm g}=GM_{\bullet}/c^{2}$. Using the condition $r_{\rm t}\lesssim r_{\rm g}$, one can infer the minimal SMBH mass, $M_{\bullet}\approx[c^{6}R_{\star}^{3}/(G^{3}m_{\star})]^{1/2}\sim 3.2\times 10^{8}\,(R_{\star}/1\,R_{\odot})^{3/2}(m_{\star}/1\,M_{\odot})^{-1/2}M_{\odot}$, for which a Solar-type star does not get disrupted but it is instead swallowed as a whole (this order-of-magnitude estimate assumes that the SMBH horizon is located on the scale of a gravitational radius). The TDE is characterized by an increased accretion of a significant fraction of the material from the shredded star, which leads to a luminosity outburst by several orders of magnitude with respect to the quiescent state. Approximately one half of the material becomes ejected from the system, and the other half circularizes to form an accretion disk [96]. Theoretical investigation as well as MHD simulations predict that the TDE outburst light curves evolve with time as power laws, $\propto t^{-5/3}$, and the observed TDEs have shown that this is a viable model, especially further away in time from the luminosity peak [97, 98]. However, full TDEs are unpredictable and require rapid follow-ups to constrain parameters of the black hole, such as the mass and the spin[99] (likely very small electric charge of SMBHs[100, 101] appears to be beyond the possibilities of current instruments, with the potential exception of the Galactic center SMBH where an upper limit was set[102]). In doing so we may learn more about the properties of black holes at the population level. More specifically, high-cadence two-band follow-up observations of TDEs are likely to shed light on the following,
@@
  • accretion flow state transitions for a sample of TDEs[105, 106], especially in synergy with X-ray monitoring. As the relative accretion rate transitions from the one close to the Eddington rate to the sub-Eddington one, the system is expected to transition from high/soft state dominated by the softer UV/optical emission ($<2$ keV) of an accretion disk to low/hard state dominated by the harder ($>2$ keV) non-thermal power-law emission of hot corona and the inner ADAF;
  • time-resolved spectral energy distributions, including radio, optical, UV, and X-ray data, of jetted TDEs, which correspond to $\sim 1\%$ of all TDEs[107]. This would allow the detailed understanding of the jet formation and collimation shortly after the TDE when the disk is still geometrically thick due to the super-Eddington accretion rate (slim disks). In case the jet is oriented close to the line of sight, the emission is enhanced due to Doppler-beaming and thus can be detected even for high-redshift ($z\gtrsim 1$) sources at cosmological distances;
@@
Figure 7: Different scenarios for the production of UV, optical, and X-ray emission in TDEs. Top left panel: Shock-at-apocenter scenario when the TDE stream collides with itself, gets shocked, and produces UV/optical emission. Fluctuations due to steam-stream collisions propagate inwards on the free-fall timescale. UV/optical light curve variability leads the X-ray emission. Top right panel: UV/optical emission is produced in the series of discrete stream-stream collisions that propagate towards the inner flow on the local viscous timescale. The X-ray emission follows the UV/optical emission. Bottom left panel: UV, optical, and the X-ray emission correspond to different radii of the elliptical accretion disk. Fluctuations propagate inwards on the local viscous timescale, which is significantly shorter than for the circular disk of a similar length-scale. In this scenario, the UV/optical variability leads the X-ray light curve. Bottom right panel: Reprocessing scenario, in which the variable X-ray emission generated within the inner accretion disk is reprocessed within the reprocessing layer at larger radii and re-radiated as a UV/optical emission. In this model, the X-ray emission leads the UV/optical emission.
@@
AGN are known to evolve with time, entering varying states, some of which are dormant. The transition between states is relatively short lived over cosmic time scales. However, changing-look AGN (CL-AGN) are likely this transitionary period and their study offers us a rare look into this process [114]. One such object is NGC 5548, known also as the “BLR on holiday”, which shows the BLR can disappear and reappear over $\sim$180 day time scale [115]. This has been used to support a dense out-flowing cloud which obstructs the BLR cloud [116]. QUVIK offers the ability to regularly monitor these objects in the UV bands and can help to find and probe these rare transitions.
@@
Another relatively rare transient event in strongly lensed AGN [117] is a microlensing HME, where a micro-caustic from the lensing galaxy tomographically scans the AGN over month-to-year time-scales [118]. Depending on the lensing configuration, micro-caustics from a lensing galaxy may amplify a region on the order of micro- to nano-arcseconds [119]. This is an external effect to the AGN and occurs in each image individually. Therefore, by monitoring these events and comparing the flux evolution between images across various wavelengths, we have the ability to probe the structure of the AGN’s central engine [120, 121, 122]. While the microlensing event occurs over weeks or months, the most interesting period of time is the peak of the HME which only lasts on the timescale of days to weeks. Recent work has gone into predicting these events from expected data from wide-field surveys [123]. With the UV emission of the AGN expected to be relatively compact compared to the optical accretion disk or the BLR, QUVIK has the ability to rapidly monitor these HMEs as they unfold and can assist in probing the central engine of the AGN.

### v2

@@ 4.1 Reverberation mapping of accretion disks
Figure 5: Top panel: Simulated light curves of a reverberating accretion disk for $M_{\bullet}=10^{9}M_{\odot}$ as modeled in the FUV (175 nm), NUV1 (260 nm), and NUV2 (325 nm) bands. Dense light curves are plotted in faded colors, while dark points represent mock observations with a cadence of 0.5 days. Lower-left panel: Discrete correlation coefficient as a function of the time lag, $\tau$, prior to bootstrapping between the FUV and NUV2 bands. Lower-right panel: Interpolated cross-correlation coefficient as a function of $\tau$ between the FUV and NUV2 bands. Within the lower panels, the blue vertical dashed line represents the peak correlation coefficient, while the pink vertical dotted line represents the centroid measured at 80 % of the peak. The black vertical dash-doted line represents the mean time delay as calculated directly from the means of the transfer functions.
@@
With the two UV bands, QUVIK has the ability to perform reverberation mapping in UV wavelengths. The AGN’s UV emission is likely due to a relatively compact region of the accretion disk near the supermassive black hole (SMBH) [74, 75, 76]. Spectra of accretion disks are more complex than a simple black body emission of typical accretion disk models (e.g. thin disk [77, 78], slim disk [79], and advection dominated accretion flows or ADAFs [80]). Magneto-hydrodynamic (MHD) simulations tend to describe the accretion flow as composed of several gaseous components with different temperatures, densities, and geometric properties  [81, 82, 83, 84, 85]. The real accretion disk is likely a combination of many of these accretion flows at different size scales. While optical continuum reverberation mapping seems to support some aspects of these accretion flow models (albeit with $\sim$2 to 3 times larger accretion disk sizes [17, 86]), there are fewer studies of UV reverberation [74, 87, 88, 89]. The contribution of UV photometric and spectroscopic observations to the multiwavelength monitoring of AGN has been essential in constraining accretion-disk temperature profiles, sizes, and the geometry and the kinematics of the circumnuclear medium; see e.g. the Swift UVOT contribution discussed in Section 1, the AGN STORM campaign on NGC 5548[90], and the AGN STORM 2 campaign on Mrk 817[91].
QUVIK will help to fill the gap by enabling the observations of a larger sample of sources more easily, especially by employing the higher cadence required to probe the smaller size scales of the inner accretion disk. In Section 5, we describe how this can be joined with other surveys at different wavelengths (optical and X-ray) to provide reverberation mapping across a wider spectrum.
@@
Here, we perform simulations of AGN light curves in order to test the efficacy of QUVIK on reverberation mapping utilizing modular code Amoeba [92]. The UV emission of the AGN is simulated as the reprocessing of a stochastic driving signal within the lamppost model, where a point-like corona irradiates the accretion disk from a position directly above the black hole [93, 94]; see also Fig. 4 for the illustration of the AGN setup that is inspired by the unified AGN model[95]. We generate light curves assuming a Schwarzschild SMBH with masses of $M_{\bullet}=10^{8},10^{9}M_{\odot}$, accreting with an Eddington ratio of 10 % and placed at a redshift of 0.03. The corona height is taken to be 6 $r_{\rm{g}}$, which is equal to the inner radius of the accretion disk, and it is assumed that it radiates with an average power of 10 % of the bolometric luminosity. We then model the driving signal as a broken power-law (BPL) function, which has a characteristic time scale ($T_{B}$) based off of equation $\log T_{B}=2.1\log(M_{\bullet}/10^{6}M_{\odot})-0.98\log(L_{\mathrm{Bol}}/10^{44}\mathrm{ergs~s}^{-1})-2.32$ [96], emphasizing that this driving signal is not an observable quantity. The power spectral density (PSD) is taken to be $\propto f^{-1}$ for low frequencies, and $\propto f^{-2}$ for high frequencies. This corresponds to a break-point timescale of 86 and 7 days for $M_{\bullet}=10^{9}$ and $10^{8}M_{\odot}$, respectively.
@@
From this simulation, the UV signals are computed assuming that the accretion disk thermally responds when the driving signal is absorbed. We computed this response by calculating the transfer functions of the accretion disk when viewed at 20[94]. These transfer functions encode the geometry and physics of the models used. It is known that inclined accretion disks do not have a significant change in their mean response function, but can change the skew of the transfer function. The UV signals are then the convolution of these transfer functions and the driving signal, to which we add 2 % Poisson noise and sample the dense light curves based on various observation strategies. We perform simulations for 3 UV bands: FUV, NUV1, and NUV2 using effective wavelengths of 175, 260, and 325 nm, respectively, where FUV and NUV2 correspond to the currently preferred QUVIK bands and NUV1 corresponds to the ULTRASAT band.
@@
To compute the time lag between the UV light curves, we then use two methods known as the Discrete Correlation Function (DCF) [97] and the Interpolated Cross-Correlation Function (ICCF) [98, 99]. These methods compute the correlation coefficient based on the time delay between two signals, where 1.0 represents perfect correlation. The time lag between two signals may be taken to be the time lag associated with the peak correlation coefficient, or alternatively, the centroid value at a specified level [93]. Common thresholds for the centroid value are 50 or 80 % of the peak correlation coefficient, which is found to give more reliable time lag measurements with respect to the peak [100, 101]. In doing so, we are less biased towards how sharp or skewed the transfer functions are. The uncertainties in the time lag are then computed using a bootstrapping method for the DCF and a Markov Chain Monte Carlo method for the ICCF, which improves our understanding of the time lag between the UV bands. For these mock observations, we set the UV light curve lengths to 180-day intervals.
While this model setup is very specific and focused on nearby, bright sources, its aim is to qualitatively explore the parameter space of possible cadences to assess the feasibility of continuum reverberation mapping in the UV domain, i.e. the measurement of the time lag between the FUV and NUV bands. In this regard, 2% uncertainty is also fixed and assumes that the AGN sources are detected with QUVIK at the signal-to-noise ratio of 100 and higher. At the moment, we do not consider a specific QUVIK UV band throughput, which will be available in the future. In addition, we also do not consider intermediate-redshift sources, for which the intrinsic rest-frame time lag $\tau_{0}$ is noticeably prolonged by the factor $(1+z)$ in the observer’s frame, which is negligible for the sources at $z<0.1$ considered here ($z=0.03$). The focus on nearby bright sources stems from the projected sensitivity of QUVIK FUV and NUV instruments, down to $m_{\rm AB}\sim 21-22$ mag, see also Fig. 3 for the analysis of the dependency of the apparent brightness on the Type I AGN redshift and the SMBH mass.
@@
QUVIK observational strategy: For the reverberation mapping in the UV bands, QUVIK can aim for AGN sources at the redshift of $z<1$, see also Fig. 3. The higher the redshift of the source, the more blueward part of the spectrum in the rest frame QUVIK can probe. In terms of a number of available sources, within $z<0.5$, there are $\sim 151$ sources with a limiting magnitude of 17 in the SDSS $u$ band (Data Release 7)[5]. For these sources, the total integration time with QUVIK to reach the signal-to-noise ratio of 100 is $\sim 100$ seconds[5]. Based on the analysis of mock observations with varied cadence and light-curve length, Zajaček et al.[5] concluded that the AGN with lower SMBH masses ($\sim 10^{7}\,M_{\odot}$) need to be monitored with a high cadence of $\sim 0.1$ days but a relatively short duration of the order of $\sim 10-100$ days is overall sufficient to capture the significant time delay between FUV and NUV bands, depending on the SMBH mass. From the past monitoring campaigns of nearby AGN, a monitoring lasting for about half a year is expected to be sufficient (see e.g. the monitoring of NGC5548 [17] with $M_{\bullet}=5\times 10^{7}\,M_{\odot}$; Mrk 817[91] with $M_{\bullet}=3\times 10^{7}\,M_{\odot}$), while the duration can decrease to a few 10 days for low SMBH-mass systems (see e.g. NGC4593[19] with $M_{\bullet}=8\times 10^{6}\,M_{\odot}$, for which the campaign lasted 26 days). On the other hand, systems with larger SMBH masses ($\sim 10^{8}\,M_{\odot}$) can be monitored with a lower cadence of $\sim 1$ day, but a longer monitoring time of at least $\sim 180$ days is required. This is consistent with the analysis presented here where the light-curve length was fixed to 180 days. However, in case more UV bands are involved in the campaign (FUV, NUV1, and NUV2), a higher-cadence monitoring at $\sim 0.1$ days is needed to detect significant time lags between the individual UV bands (FUV-NUV1 vs. FUV-NUV2).
@@ 4.2 Stochastic nuclear transients
where $\kappa$ is of the order of unity and depends on the radial density profile of the star. Eq. (3) expresses the tidal radius conveniently in gravitational radii, $r_{\rm g}=GM_{\bullet}/c^{2}$. Using the condition $r_{\rm t}\lesssim r_{\rm g}$, one can infer the minimal SMBH mass, $M_{\bullet}\approx[c^{6}R_{\star}^{3}/(G^{3}m_{\star})]^{1/2}\sim 3.2\times 10^{8}\,(R_{\star}/1\,R_{\odot})^{3/2}(m_{\star}/1\,M_{\odot})^{-1/2}M_{\odot}$, for which a Solar-type star does not get disrupted but it is instead swallowed as a whole (this order-of-magnitude estimate assumes that the SMBH horizon is located on the scale of a gravitational radius). The TDE is characterized by an increased accretion of a significant fraction of the material from the shredded star, which leads to a luminosity outburst by several orders of magnitude with respect to the quiescent state. Approximately one half of the material becomes ejected from the system, and the other half circularizes to form an accretion disk [104]. Theoretical investigation as well as MHD simulations predict that the post-TDE accretion rate evolves with time as a power law, $\propto t^{-5/3}$, and the observed TDEs have shown that this is a viable model for the luminosity evolution as well, especially further away in time from the luminosity peak [105, 106]. However, full TDEs are unpredictable and require rapid follow-ups to constrain parameters of the black hole, such as the mass and the spin[107] (likely very small electric charge of SMBHs[108, 109] appears to be beyond the possibilities of current instruments, with the potential exception of the Galactic center SMBH where weak charge constraints can be set[110, 111]). In doing so we may learn more about the properties of black holes at the population level. More specifically, high-cadence two-band follow-up observations of TDEs are likely to shed light on the following,
@@
  • accretion flow state transitions for a sample of TDEs[114, 115], especially in synergy with X-ray monitoring. As the relative accretion rate transitions from the one close to the Eddington rate to the sub-Eddington one, the system is expected to transition from high/soft state dominated by the softer UV/optical emission (<2<2 keV) of an accretion disk to low/hard state dominated by the harder (>2>2 keV) non-thermal power-law emission of hot corona and the inner ADAF;
  • time-resolved spectral energy distributions, including radio, optical, UV, and X-ray data, of jetted TDEs, which correspond to $\sim 1\%$ of all TDEs[116]. This would allow the detailed understanding of the jet formation and collimation shortly after the TDE when the disk is still geometrically thick due to the super-Eddington accretion rate (slim disks). In case the jet is oriented close to the line of sight, the emission is enhanced due to Doppler-beaming[117, 118] and thus can be detected even for high-redshift ($z\gtrsim 1$) sources at cosmological distances;
@@
Figure 7: Different scenarios for the production of UV, optical, and X-ray emission in TDEs. a) Shock-at-apocenter scenario when the TDE stream collides with itself, gets shocked, and produces UV/optical emission. Fluctuations due to steam-stream collisions propagate inwards on the free-fall timescale. UV/optical light curve variability leads the X-ray emission. b) UV/optical emission is produced in the series of discrete stream-stream collisions that propagate towards the inner flow on the local viscous timescale. The X-ray emission follows the UV/optical emission. c) UV, optical, and the X-ray emission correspond to different radii of the elliptical accretion disk. Fluctuations propagate inwards on the local viscous timescale, which is significantly shorter than for the circular disk of a similar length-scale. In this scenario, the UV/optical variability leads the X-ray light curve. d) Reprocessing scenario, in which the variable X-ray emission generated within the inner accretion disk is reprocessed within the reprocessing layer at larger radii and re-radiated as a UV/optical emission. In this model, the X-ray emission leads the UV/optical emission.
@@
AGN are known to evolve with time, entering varying states, some of which are dormant. The transition between states is relatively short lived over cosmic time scales. However, changing-look AGN (CL-AGN) are likely this transitionary period and their study offers us a rare look into this process [125]. CL-AGN seem to be associated with the rapid changes in the structure of the innermost accretion flow, e.g. the case of 1ES 1927+654 seems to indicate the destruction of the corona and its recreation on a yearly timescale[126, 127]. A possible interpretation of such abrupt changes involves the interaction of the denser circumnuclear gaseous material (e.g. post-TDE streams) with the accretion flow, which initially enhances the accretion rate and subsequently causes the depletion of the inner accretion flow and the destruction of hot corona. After some time, the inner disk is refilled and the corona reforms as well. Supermassive binary black holes shortly before the merger are also candidates for curious and specific variability patterns. While acting as relatively normal AGN before the merger, where the emission originates in the circumbinary disk and minidisks around each black hole, shortly after the merger the X-ray emission nearly disappears while the UV emission originating in the circumbinary disk is largely unaffecred. This could be a tell-tale sign of supermassive binary black holes merging[128]. QUVIK offers the ability to regularly monitor these objects in the UV bands and can help to find and probe these rare transitions.
Other variability phenomena are associated with the changes in the variability pattern rather than the “look” itself. Such a special case is NGC 5548, which exhibited the phenomenon known as the “BLR on holiday”, which shows that the BLR emission variability can decouple from the photoionizing continuum variability over $\sim$180 day time scale [129]. It was suggested that the “holiday” was caused by a dense out-flowing cloud which obscured the underlying continuum, breaking the correlation with the BLR emission [130]. In fact, the disk winds with a wide range of column densities may be a ubiquitous phenomenon. When they are fully transparent, standard BLR reverberation takes place. When the density increases above the threshold value, it induces the decoupling between the continuum and the BLR variability, which can address “BLR holidays”. For even larger densities over a certain time span, i.e. when the disc wind prevents the continuum emission from reaching BLR clouds at larger distances, broad lines can disappear as seen in some CL-AGN [130].
@@
Another relatively rare transient event in strongly lensed AGN [131] is a microlensing high-magnification event (HME), where a micro-caustic from the lensing galaxy tomographically scans the AGN over month-to-year time-scales [132]. Depending on the lensing configuration, micro-caustics from a lensing galaxy may amplify a region on the order of micro- to nano-arcseconds [133]. This is an external effect to the AGN and occurs in each image individually. Therefore, by monitoring these events and comparing the flux evolution between images across various wavelengths, we have the ability to probe the structure of the AGN’s central engine [134, 135, 136]. While the microlensing event occurs over weeks or months, the most interesting period of time is the peak of the HME which only lasts on the timescale of days to weeks. Recent work has gone into predicting these events from expected data from wide-field surveys [137]. With the UV emission of the AGN expected to be relatively compact compared to the optical accretion disk or the BLR, QUVIK has the ability to rapidly monitor these HMEs as they unfold and can assist in probing the central engine of the AGN.
QUVIK observational strategy: In order to distinguish among different origins of the UV light in TDE sources, see Fig. 7, or to study processes in CL-AGN well-sampled UV, optical, and X-ray light curves need to be compared with model light curves at the corresponding wave-bands. Detailed modeling of such processes is beyond the scope of this overview. Here we just compare relevant timescales that the UV mission should be able to resolve for this purpose, see Table 1. The shortest timescales of a fraction of the day are dynamical and light-crossing timescales. The thermal timescale of a few days is related to the heating/cooling-front propagation. The longest timescales are related to the perturbation propagation and viscous processes ($\gtrsim$ year). In general, since all of the timescales are proportional to the SMBH mass, they can vary by orders of magnitude. In addition, the viscous timescale, on which accretion processes take place, can be significantly shortened if instead of the whole disc at the distance $r$ only its narrow ring of the width $\Delta r$ is affected by e.g. an instability. Then the viscous timescale is $\tau_{\rm visc}(\Delta r)\approx(\Delta r)/r)\tau_{\rm visc}$ [138]. Therefore an initial high-cadence monitoring of any nuclear transient at $\sim 0.1$ day can resolve several dynamical processes taking place in the inner accretion disk or affecting its narrower part. From an observational point of view, the duration of the monitoring as well as the cadence will need to be adjusted for intermediate-redshift sources; e.g. for $z\sim 1$, the necessary duration of the campaign would be prolonged by $(1+z)\sim 2$, while the characteristic variability timescale would also get stretched by the same factor. However, for intermediate-redshift sources, QUVIK can probe the progressively shorter-wavelength spectral regions towards the FUV domain in the rest frame of the source ($175\,{\rm nm}/(1+z)\sim 87.5\,{\rm nm}$ in the FUV band and $325\,{\rm nm}/(1+z)\sim 162.5\,{\rm nm}$ in the NUV band), for which the variability timescale tends to be smaller, and therefore the monitoring cadence should remain approximately the same to capture it. In order to observe the UV transients at the satisfactory signal-to-noise ratio, such as TDEs, the redshift distribution of the detectable sources peaks towards low-redshift, nearby nuclei. Assuming the SMBH accreting at the Eddington limit shortly after the star is tidally disrupted in its vicinity, the bolometric luminosity is $L_{\rm bol}\sim 1.26\times 10^{44}\,{\rm erg\,s^{-1}}$ for $M_{\bullet}=10^{6}\,M_{\odot}$. Most of this luminosity is initially radiated away in the optical/UV domain, hence $L_{\rm UV}\sim 10^{44}\,{\rm erg\,s^{-1}}$ at $\lambda=300$ nm. This yields the apparent AB magnitude of $m_{\rm AB}=-2.5\log{F_{\nu}}-48.60\sim 17.4$ for $z=0.1$, which is well within the QUVIK sensitivity range. Adopting the conservative sensitivity limit $m_{\rm AB}\lesssim 21.5$ mag for QUVIK, we obtain the limiting redshift of $z\lesssim 0.53$ for the detectability of bright UV transients (TDEs). For this redshift, the corresponding time dilation factor is $\sim 1.5$, while at the same time QUVIK will probe the spectral region at $325\,{\rm nm}/(1+z)\sim 217\,{\rm nm}$ in the rest frame of the source.
Table 1: Comparison of relevant timescales for the SMBH accretion disks. $M_{\bullet}$ stands for the SMBH mass, $r$ for the distance from the SMBH, $\alpha$ is a viscous parameter, and $h/r$ is a scale-height-to-radius ratio.
Time scale Relation
Dynamical $\tau_{\rm dyn}=12.2\left(\frac{r}{20~r_{\rm g}}\right)^{3/2}\left(\frac{M_{\bullet}}{10^{8}~M_{\odot}}\right)$ hours
Light-crossing $\tau_{\rm lc}=2.7\left(\frac{r}{20~r_{\rm g}}\right)\left(\frac{M_{\bullet}}{10^{8}~M_{\odot}}\right)$ hours
Thermal $\tau_{\rm th}=5.1\left(\frac{\alpha}{0.1}\right)^{-1}\left(\frac{r}{20~r_{\rm g}}\right)^{3/2}\left(\frac{M_{\bullet}}{10^{8}~M_{\odot}}\right)$ days
Perturbation propagation $\tau_{\rm prop}=509\left(\frac{\alpha}{0.1}\right)^{-1}\left(\frac{h/r}{0.01}\right)^{-1}\left(\frac{r}{20~r_{\rm g}}\right)^{3/2}\left(\frac{M_{\bullet}}{10^{8}~M_{\odot}}\right)$ days
Viscous $\tau_{\rm visc}=139\left(\frac{\alpha}{0.1}\right)^{-1}\left(\frac{h/r}{0.01}\right)^{-2}\left(\frac{r}{20~r_{\rm g}}\right)^{3/2}\left(\frac{M_{\bullet}}{10^{8}~M_{\odot}}\right)$ years